b) Jack also wants to cover his backyard with grass. Grass costs $2 per square metre, how much will it cost to fill the backyard. (2 marks)









3. Jack wants to fence his rectangular backyard. The backyard is 20 metres long and 40 metres wide. The cost of materials for the fence is $10 per metre.
a) How much is the total cost of putting the fence around the backyard. (2 marks)










b) Jack also wants to cover his backyard with grass. Grass costs $2 per square metre, how much will it cost to fill the backyard. (2 marks)
1
Expert's answer
2020-05-28T19:41:46-0400

a) What is the perimeter of the backyard?


"P=2(20+40)=120 (m)"

How much is the total cost of putting the fence around the backyard?


"10\cdot120=1200(\$)"

The total cost of putting the fence around the backyard is $1200.


b) What is the area of the backyard?


"A=20(40)=800 (m^2)"

How much will it cost to fill the backyard?


"2\cdot800=1600(\$)"

It will cost $1600 to fill the backyard.